A stock solution of copper, 1000.0 (± 0.3) mg Cu /L, is diluted tenfold by diluting 10.00 (± 0.04) mL stock to a final volume of 100.00 (± 0.05) mL. The concentration of the diluted solution, calculated using the formula Cstock.Vstock = Cdiluted.Vdiluted, is of course 100 mg Cu /L.

Cdiluted = 1000.0(±0.3) x 10.00(±0.04) / 100.00(±0.05)

Calculate the % relative uncertainty in the concentration of the 100 mg/L diluted solution. The numbers in parenthesis are absolute uncertainties. Enter your answer to two significant figures in the format 0.0E0. Do NOT include the % sign.
1
Expert's answer
2020-10-14T05:47:35-0400

Milli-moles of KI=50×0.04=2Milli-equivalents of Na2​S2​O3​ reacted with 10 mL solution=20×2.5×10−3=50×10−3
Milli-equivalent of I− left unreacted in 100 mL solution=50×10−3×10
100​=0.5Milli-moles of I− reacted with CuCO3​=2−0.5=1.5∵2Cu2++5I−→2CuI+I3−​Milli-moles of Cu2+=5
2​×1.5=0.6Weight of CuCO3​=0.6×10−3×123.5=0.0741Weight % of CuO3​=0.1
0.0741​×100=74.1